The number of calls received by ABC Company follows Poisson distribution with an average of 6 calls per hour.

(a) Find the probability that there will be exactly 3 calls received in an hour. Answer: 

(b) Find the probability that there will be at least 2 calls received in an hour. Answer: 

(c) Find the probability that there will be exactly 2 calls received in 10 minutes. Answer: 



1
Expert's answer
2021-06-14T13:07:19-0400

"(a) \\
Solution:\\
P(3)\\
\space Probability\space of\space exactly\space 3\space occurrences\\


mean=λ=6\\
\space and\space \\
x=3\\
\space \space \space apply\space the\space poisson\space probability\space formula:\\
P
(
x
)
=
\frac{e^{
−
λ}.
⋅
λ^
x}
{x
!}\\
where\space \\
x
\space is\space the\space number\space of\space occurrences,\space \\
λ
\space is\space the\space mean\space number\space of\space occurrences,\space and\space \\
e
\space is\space the\space constant\space 2.718.\\\space Substituting\space in\space values\space for\space this\space problem,\space \\
x
=
3\\
\space and\space \\
λ
=
6\\
,\space we\space have\\
P
(
x
)
=
\frac{e^{
−
6}.
⋅
6^
3}
{3
!}\\
Evaluating\space the\space expression,\space we\space have\\
P
(
3
)
=
0.089235078359989\\
-------------------------\\
probability \space \space that \space there \space will \space be \space at \space least \space 2 \space calls \space received \space in \space an \space hour. \space \\
P(x \space \geq \space 2)= \space 1-P(x \space <2) \space \\
Now \space we \space find \space P(x<2)\\ \space 
P(x<2)=P(x \space =0) \space + \space P(x \space =1) \space \\

=
\frac{e^{
−
6}.
⋅
6^
0}
{0
!}+
\frac{e^{
−
6}.
⋅
6^
1}
{1
!}\\
Evaluating\space \space the\space \space expression,\space \space we\space \space have\\
P
(
0
)+P(1)
=
0.00247875218+0.0148725131\\
=0.01735126528\\
P(x \space \geq \space 2)= \space 1-0.01735126528 \space \\
=0.98264873472 \\
---------------------\\
(C) \space \\
probability \space that \space there \space will \space be \space exactly \space 2 \space calls \space received \space in \space 10 \space minutes \space \\
Given \space 6 \space phone \space call \space =60 \space min\\
1 \space phone \space call \space =10 \space min\\
P(x=2)=\frac{e^{(-1) \space } \space 1^2 \space }{2!}\\
\\
0.183939721"